Let's say that the Federal Reserve purchases $1 Million worth of U.S. Treasury bonds from a bond dealer in the open market, and the dealer's bank credits the dealer's account. The required reserve ratio is 15 percent, and the bank typically lends any excess reserves immediately.

a) Assuming that no currency leakage occurs, how much will the bank be able to lend to its customers following the Fed's purchase? Please explain and show your calculations.

b) Using the simple money multiplier, how much money will be created (assuming every dollar is lent out)? Please explain and show your calculations.
Answer #1


A)

Total deposits with bank = $1 million

Required reserves = 15% of $1 = $0.15 million

Excess reserves = $1 - $0.15 = $0.85 million

Thus, the bank will be able to lend $0.85 million to its customers

B)

Money multiplier = 1/rr = 1/0.15 = 6.67

Total increase in money supply = $1 x 6.67 = $6.67 million

Thus, money supply will increase by $6.67 million